# Detection of Non-recorded Word Senses in English and Swedish

### v1

@@ Results.
When optimizing for $F_{0.3}$-scores, model E4_COS produces the best results for English in the cross validation (see Appendix D for an overview and Appendix E for an example of an individual evaluation round). However, since the value deviates only minimally from the performance of E4_SUB_SPR, we decide to use the latter, as the standard deviation of the scores over ten rounds of cross-validation is lower. For Swedish, we decide on G3_COS, as it achieves the best results. All average model performances exceed the baselines.
              Phase I: Random sample                 Phase II: Model prediction
English Swedish English Swedish
Instances 1165 1202 1208 1400
Usages 474 706 322 1001
Label distribution (0, 1, -) (1840, 1651, 4) (1294, 2104, 208) (2151, 1462, 11) (2529, 1218, 456)
Excluded instances 2 87 5 109
Remaining usages 473 674 322 927
Assigned 428 562 277 327
Unassigned 45 (9.5%) 95 (16.6%) 45 (13.98%) 600 (64.725%)
Table 5: Statistics of the first (left) and second (right) human annotation phases
@@ 6.2 Phase II: Model prediction
Then, usages for evaluation are chosen from the top of this list. At most eight usages for the same headword are sampled to ensure a broad analysis. These usages are combined with all eligible senses from the respective dictionary. The finally resulting sample has roughly the same size as in the first annotation phase. This is uploaded to PhiTag and annotated in parallel to the first phase. The annotation analysis is described above in Section 5.2.
@@ Results.
Table 5 summarizes the annotation results from both phases. The second (prediction) phase yields diverse results. While both models successfully increase the number of unassigned usages in the sample data compared to the first phase (random sample), the Swedish model is considerably more successful: Out of 1000 predictions, the Swedish model is correct in almost two thirds of cases according to human judgment. In the predictions of the Swedish model, there is a significantly higher number of usages with few senses (1,400 instances for 1,001 usages vs. 1,208 instances for 322 usages, even though we have similar numbers of senses per headword for both dictionaries, as shown in Table 3).
@@
Differentiating between modern and historical instances indicates a commonality between both languages (see Appendix G): In the first phase, unassigned senses were almost equally distributed between modern and historical usages in the English data. In the second phase, however, the proportion of correctly predicted historical usages hardly differs from the random sample of the first phase while the correct predictions of modern usages increases by two thirds. We observe a similar trend in Swedish: Although the distribution of unassigned usages is very different between modern and historical in the first Swedish phase with proportionately almost twice as many unassigned usages in the historical data, a stronger relative increase in the modern usages can also be observed here.
@@
A partial manual analysis carried out on the Swedish predictions yields very similar conclusions. We find several non-recorded senses such as the following usages of avtryck:
  • usage: Men framförallt är det Katarinas beslutsamhet och kompetens som gett avtryck.
    “But above all, it is Katarina’s determination and competence that have left an impression.”
  • usage: Kanadas premiärminister Justin Trudeau säger i ett tårfyllt tal att han minns sitt första möte med drottningen som barn och att hon för alltid kommer att ha lämnat ett avtryck på landet.
    “Canadian Prime Minister Justin Trudeau says in a tearful speech that he remembers his first meeting with the Queen as a child and that she will forever have left an imprint on the country.”
  • senses: ‘efterlämnat spår av föremål som tryckts mot ett mjukare underlag’: “traces of objects that have pressed against a softer surface”; ‘(resultat av) avdrag av plåt eller tryckform’: “(result of) impression of plate or printing form”
Further examples, amongst others, include usages of avslut, kanta, internalisera, kidnappa, organisk, partnerskap and pilot. However, similar to English, true positives contain many problematic cases: Besides multi-word expressions, proper names and lemmatization are major problems: For instance, for the noun ode we detected the usage På pizzeria Oden (“At the pizzeria Oden …”). And, the word sabbat can be a noun (Sabbath) or a form of the colloquial verb sabba (“destroy”). We also observe a number of cases where our extraction of the dictionary was not complete and thus the predicted sense was actually recorded in SO, though not being present in our data (see Section 8).
In summary, we found a number of non-recorded senses for both languages. Our predictions for Swedish will be included in SO in the future. Most errors amongst the true positives concern the matching of headwords to word usages. As we used very simple heuristics to sample usages, we see possibility for improvement of the model.
@@ 8 Limitations
Due to the exclusion of non-primary synsets in the first annotation phase, an unrestricted comparison between the two English annotation phases is not possible. To overcome this limitation, future experiments ought to ensure that pre- and post-prediction annotation conditions closely mirror each other.
Moreover, the baselines described in Section 6.1 were run independently for each model. Future work should fix these across models in order to make model performances more comparable. Future work should also investigate the use of ensemble methods, or combinations that use both gloss and example representations (e.g., by averaging vectors).
@@
Another limitation of our work is the treatment of SO: In our extraction from the SO database, [^15] The code is available at https://github.com/ChangeIsKey/SO-extract-db. A few sense glosses were extracted with errors. we did not consider sub-senses of headwords (e.g. figurative uses) and only kept a word’s main senses. This led to an over-classification for non-recorded senses.
@@ Acknowledgments
This paper is a revised version of Lautenschlager (2023) on which we carried out several additional analyses. Dominik Schlechtweg and Emma Sköldberg have been funded by the research program ‘Change is Key!’ supported by Riksbankens Jubileumsfond (under reference number M21-0021) during the creation of this study. We thank Haim Dubossarsky and Nina Tahmasebi for contributions to this project.
Appendix A Example of an annotation instance
Example of an annotation instance for a usage of the word relative. The word itself is highlighted, one gloss is marked for annotation and all glosses are given as additional information.
Appendix B Wordnet in a dictionary form
@@
Appendix C Selection of evaluation metric
@@
Appendix D Cross-validation overview
@@
Appendix E Cross-validation in round 10 of model E4_COS for English
@@
Appendix F Inter-annotator agreement
English Swedish
All Modern Historic All Modern Historic
A1 vs. A2 0.331 0.392 0.276 0.408 0.506 0.317
A1 vs. A3 0.273 0.333 0.220 0.575 0.614 0.527
A2 vs. A3 0.561 0.580 0.544 0.524 0.655 0.353
Full 0.389 0.434 0.349 0.480 0.588 0.371
English Swedish
All Modern Historic All Modern Historic
A1 vs. A2 0.335 0.310 0.347 0.539 0.566 0.456
A1 vs. A3 0.342 0.291 0.397 0.558 0.546 0.552
A2 vs. A3 0.464 0.490 0.374 0.608 0.641 0.526
Full 0.384 0.368 0.375 0.559 0.584 0.490
Appendix G Comparison of annotation results per corpus

### v2

@@ Results.
When optimizing for $F_{0.3}$-scores, model E4_COS produces the best results for English in the cross validation (see Appendix G for an overview and Appendix H for an example of an individual evaluation round). However, since the value deviates only minimally from the performance of E4_SUB_SPR, we decide to use the latter, as the standard deviation of the scores over ten rounds of cross-validation is lower. For Swedish, we decide on G3_COS, as it achieves the best results. All average model performances exceed the baselines.
English Swedish
Instances 1165 1202
Usages 474 706
Label dist. (0, 1, -) (1840, 1651, 4) (1294, 2104, 208)
Excluded instances 2 87
Remaining usages 473 674
Assigned 428 562
Unassigned 45 (9.5%) 95 (16.6%)
Table 5: Statistics of annotation phase I (random sample)
English Swedish
Instances 1208 1400
Usages 322 1001
Label dist. (0, 1, -) (2151, 1462, 11) (2529, 1218, 456)
Excluded instances 5 109
Remaining usages 322 927
Assigned 277 327
Unassigned 45 (13.98%) 600 (64.725%)
Unassigned baseline 45 (9.5%) 95 (16.6%)
Table 6: Statistics of annotation phase II (model prediction). The bottom row gives the annotation result from phase I obtained from randomly sampling usages as a baseline.
@@ 6.2 Phase II: Model prediction
Then, usages for evaluation are chosen from the top of this list. At most eight usages for the same headword are sampled to ensure a broad analysis. These usages are combined with all eligible senses from the respective dictionary. The finally resulting sample has roughly the same size as in the first annotation phase. This is annotated in parallel to the first phase. The annotation analysis is described above in Section 5.2.
@@ Results.
Table 6 summarizes the annotation results from phase II, yielding diverse results. While both models successfully increase the number of unassigned usages in the sample data compared to the first phase (random sample, see column “Unassigned baseline”), the Swedish model is considerably more successful: Out of 1000 predictions, the Swedish model is correct in almost two thirds of cases according to human judgment. In the predictions of the Swedish model, there is a significantly higher number of usages with few senses (1,400 instances for 1,001 usages vs. 1,208 instances for 322 usages, even though we have similar numbers of senses per headword for both dictionaries, as shown in Table 3).
@@
Differentiating between modern and historical instances indicates a commonality between both languages (see Appendix I): In the first phase, unassigned senses were almost equally distributed between modern and historical usages in the English data. In the second phase, however, the proportion of correctly predicted historical usages hardly differs from the random sample of the first phase while the correct predictions of modern usages increases by two thirds. We observe a similar trend in Swedish: Although the distribution of unassigned usages is very different between modern and historical in the first Swedish phase with proportionately almost twice as many unassigned usages in the historical data, a stronger relative increase in the modern usages can also be observed here.
@@ 8 Limitations
We only realised throughout the experiments for phase I that we had missed a number of senses that were recorded as secondary synsets. Hence, we decided to include these for phase II. Due to the exclusion of non-primary synsets in the first annotation phase, an unrestricted comparison between the two English annotation phases is not possible. To overcome this limitation, future experiments ought to ensure that pre- and post-prediction annotation conditions closely mirror each other.
Moreover, and the baselines described in Section 6.1 were run independently for each model. Future work should fix these across models in order to make model performances more comparable. Future work should also investigate the use of ensemble methods, or combinations that use both gloss and example representations (e.g., by averaging vectors).
@@
Another limitation of our work is the treatment of SO: In our extraction from the SO database, we did not consider sub-senses of headwords (e.g. figurative uses) and only kept a word’s main senses. This led to an over-classification for non-recorded senses. [^16] The code is available at https://github.com/ChangeIsKey/SO-extract-db. A handful of sense glosses were extracted with errors.
@@ Acknowledgments
This paper is a revised version of Lautenschlager (2023) on which we carried out several additional analyses. Dominik Schlechtweg and Emma Sköldberg have been funded by the research program ‘Change is Key!’ supported by Riksbankens Jubileumsfond (under reference number M21-0021) during the creation of this study. We thank Haim Dubossarsky and Nina Tahmasebi for contributions to this project. This work has been made possible by code created by iguanodon.ai.
Appendix A Wordnet in a dictionary form
@@
Appendix B Gloss and secondary gloss
Find below the entry for the Swedish noun svindel. The word has two main senses, listed in definitions. While the first sense is only described by a gloss, the second sense has both a main gloss (“(ekonomiskt) bedrägeri i större skala") and a secondary gloss (“i större skala").
        "key": 145357,
        "word": "svindel",
        "nature": "subst.",
        "definitions": [
                "gloss": "+yrsel som uppkommer vid vistelse p hga hjder",
                "sub_gloss": "",
                "sub_entries": [
                        "gloss": "v. om likn. knsla som uppsttt av annan orsak",
                        "sub_gloss": "",
                        "sub_entries": [],
                        "examples": [
                            "han knde svindel vid tanken p hur mycket pengar han hade ansvar fr"
                        "year": "1668"
                "examples": [
                    "hon fick svindel uppe i tornet"
                "year": "1668"
                "gloss": "(ekonomiskt) bedrgeri",
                "sub_gloss": "i strre skala",
                "sub_entries": [],
                "examples": [],
                "year": "1879"
Appendix C Example of an annotation instance
Example of an annotation instance for a usage of the word relative. The word itself is highlighted, one gloss is marked for annotation and all glosses are given as additional information. The annotator is asked to choose between label “1” (sense gloss fits) and label “0” (sense gloss does not fit). The annotators also always have the option to choose that no specification is possible (label “-”).
Appendix D Inter-annotator agreement
English Swedish
All Modern Historic All Modern Historic
A1 vs. A2 0.331 0.392 0.276 0.408 0.506 0.317
A1 vs. A3 0.273 0.333 0.220 0.575 0.614 0.527
A2 vs. A3 0.561 0.580 0.544 0.524 0.655 0.353
Full 0.389 0.434 0.349 0.480 0.588 0.371
English Swedish
All Modern Historic All Modern Historic
A1 vs. A2 0.335 0.310 0.347 0.539 0.566 0.456
A1 vs. A3 0.342 0.291 0.397 0.558 0.546 0.552
A2 vs. A3 0.464 0.490 0.374 0.608 0.641 0.526
Full 0.384 0.368 0.375 0.559 0.584 0.490
Appendix E Example of evaluation pipeline
Consider the annotation instance for the word relative from Appendix C. Annotators will likely assign this usage to the gloss “a person related by blood or marriage”. Let’s assume that the annotation outcome for this instance from the three annotators is $(1,1,-)$. These are aggregated by majority vote as described in Section 5.1, in this case $1$. Hence, this usage is considered assigned. For clarity, we further assume that annotations for all other glosses of this usage are aggregated as 00, i.e., the usage is unambiguously assigned to the above-mentioned sense gloss. The annotation instance is then manipulated to perform model selection during the simulation (masking) and cross-validation procedure described in Section 6.1. For each round of the simulation procedure, there are two possibilities: (i) the assigned sense is randomly masked, or (ii) not. For (i), models are optimized to predict the usage as unassigned (unknown), i.e., they should not assign this usage to any remaining (non-masked) glosses/senses of the word relative. For (ii), models are optimized to predict the usage as assigned (known), i.e., the usage should be assigned to any of the glosses/senses. (Although we know the correct assignment, we do not distinguish between possible assignments.) Gloss models (G) create their prediction for each instance by taking the word usage and the list of possible glosses as input, manipulating these according to Table 4, embedding usage and sense gloss, computing each pairwise cosine similarity and deciding on the assignment according to a threshold on the similarity. The threshold is optimized globally over all prediction instances. Example models (E) work similarly with the difference that they do not embed a sense according to their gloss, but according to observed examples of that sense.
Appendix F Selection of evaluation metric
@@
Appendix G Cross-validation overview
@@
Appendix H Cross-validation in round 10 of model E4_COS for English
@@
Appendix I Comparison of annotation results per corpus